7+2(3x-5)=8-3(2x+1)
Answer by checkley75(3666)

7+2(3X-5)=8-3(2X+1)
7+6X-10=8-6X-3
6X+6X=8-3-7+10
12X=8
X=8/12
X=2/3 ANSWER